OpenStudy (marsmcnugget):

What is the solution to the system of equations? x+y=2 x-y=4

OpenStudy (marsmcnugget):

i have to get my work done fast so i can go see my little sister i havnt seen in months i will metal and fan

pooja195 (pooja195):

x+y=2for x: x+y+−y=2+−y(Add -y to both sides) x=−y+2

OpenStudy (marsmcnugget):

and i rly suck at math

pooja195 (pooja195):

Now we know what x equals Substitute−y+2 for x in x−y=4: x−y=4 −y+2−y=4 Solve for y

pooja195 (pooja195):

If you aren't willing to learn I can't help sorry.